The floor of a rectangular storage room is divided into two right triangles by drawing a diagonal. One leg of the resulting triangle, the width of the room, is 4 feet more than two times the other leg, the length of the room. The hypotenuse is 26 feet. Find the length and the width of the room.
Answer by jhaashok630(18)

Let x = one leg length of triangle

the width of the room, is 4 feet more than two times the other leg
width = 2x+4
use pythagoreous therem
One leg = x other leg = 2x+4 hyptenuse = 26
x^2+(2x+4)^2=26^2
x=10 length = 10 feet and width = 24 feet